OpenStudy (anonymous):

I desperately need help with this question. d^2x/dt^2 + dx/dt + Sin x = 0 Find the fixed points Thanks

OpenStudy (goformit100):

Welcome to Open Study @Harrisoncd2468

OpenStudy (anonymous):

\[\begin{align*}\frac{d^2x}{dt^2}+\frac{dx}{dt}+\sin x&=0\\\\ \frac{d^2x}{dt^2}+\frac{dx}{dt}&=-\sin x\end{align*}\] The homogeneous solution: \[x''+x'=0\] gives the characteristic equation \[r^2+r=0~~\iff~~r(r+1)-0~~\iff~~r=0,~r=-1\] Solution: \(x_h=C_1e^{0x}+C_2e^{-x}=C_1+C_2e^{-x}\) The nonhomogeneous solution: Solve by the method of undetermined coefficients. Suppose \(x_p=A\sin t+B\cos t\) is a solution, then \[\begin{align*}x_p&=A\sin x+B\cos x\\ {x_p}'&=A\cos x-B\sin x\\ {x_p}''&=-A\sin x-B\cos x\end{align*}\] Substituting into the DE, you have \[\begin{align*}(-A\sin x-B\cos x)+(A\cos x-B\sin x)&=-\sin x\\ (-A-B)\sin x+(A-B)\cos x&=-\sin x \end{align*}\] which gives the system \[\begin{cases}-A-B=-1\\ A-B=0\end{cases}~~\Rightarrow~~A=B=\frac{1}{2}\] So, the general solution is \[\begin{align*}x(t)&=x_h+x_p\\ &=C_1+C_2e^{-x}+\frac{1}{2}\sin x+\frac{1}{2}\cos x \end{align*}\]

OpenStudy (anonymous):

Not too sure about this "fixed point" business. I haven't come across the term before.